# A survey on Concept-based Approaches For Model Improvement

### 21 Mar 2024 v1

@@ Abstract
The focus of recent research has shifted from merely increasing the Deep Neural Networks (DNNs) performance in various tasks to DNNs, which are more interpretable to humans. The field of eXplainable Artificial Intelligence (XAI) has observed various techniques, including saliency-based and concept-based approaches. Concept-based approaches explain the model’s decisions in simple human understandable terms called Concepts. Concepts are human interpretable units of data and are the thinking ground of humans. Explanations in terms of concepts enable detecting spurious correlations, inherent biases, or clever-hans. With the advent of concept-based explanations, there have been various concept representation methods and automatic concept discovery algorithms. Some recent methods use concepts for post-hoc model disentanglement evaluation, while others use them for ante-hoc training. The concept-based approaches are new, with many representations coming up, and there is very limited work on Concept-based Model improvement. We provide a systematic review and taxonomy of various concept representations and their discovery algorithms in DNNs, specifically in vision. We also provide details on concept-based model improvement literature, which is the first to survey concept-based model improvement methods.
@@ 1 Introduction
With the increasing use of deep learning in various tasks, there is rapid growth in the XAI literature. One of the most promising methods of XAI is concept-based approaches, which explain model predictions in human-understandable units called Concepts Kazhdan et al. (2021). Concepts are linked to their origin in NeuroScience Literature, in which they are defined as ideas derived or inferred from facts Donham (2010) and hold the human’s mental world together Murphy (2002). It has been observed that humans learn by creating concept models of things around them, helping them gather generalized knowledge Lake et al. (2017). Model explanations in terms of concepts not only lead to better interpretability but also help identify inherent biases in the model or clever-hans effects.
@@
The interpretability research aims to explain the model’s decisions in simple human understandable terms Zhang et al. (2021b). Concepts, by their definition, are humanly understandable and, hence, a good choice for interpretability. Recent XAI literature has observed a surge in concept-oriented explanations. There have been many techniques that derive concept-based explanations from pre-trained models (post-hoc), while some methods do it while training the model (ante-hoc). Kim et al. (2018) paved the way for post-hoc concept-based explanations of models using Concept Activation Vectors, while a parallel work Fong and Vedaldi (2018) explained Convolutional Neural Networks (CNN’s) filters in terms of concepts. Many works Bai et al. (2022); Chen et al. (2020); Chang (2018b); Bontempelli et al. (2023) build on the above concept representations in a post-hoc manner. Concept Bottleneck Models (CBMs) Koh et al. (2020) derive concept explanations ante-hoc, putting the weightage of concept explanations on the model while training. Recent work by Yuksekgonul et al. (2022) extended CBMs to post-hoc explanations.
@@
Concepts are essential for both human-like learning and better generalization by models. A natural intuition for human-like learning is making the model match the human attention for several tasks as done by Shen et al. (2021); Gao et al. (2022b); Liao et al. (2022); He et al. (2022); Helgstrand and Hultin (2022). One major limitation of such approaches is the requirement of additional human-level attention data. Chang (2018a) formalizes Concept Oriented Deep Learning (CODL) for human alike learning involving concepts. They argue that CODL overcomes several limitations of deep learning, involving concerns of interpretability, transferability, contextual adaptation, and humongous training data requirements. For the scope of this work, we focus on concept-based literature and refrain from providing generic definitions of interpretability/explainability (or XAI) literature, referring the readers to Zhang et al. (2021b); Du et al. (2019); Linardatos et al. (2020); Vojíř and Kliegr (2020); Akhtar (2023) which present surveys on XAI methods. Another concept-inspired learning paradigm is Neuro-Symbolic learning, which involves learning like humans using symbolic reasoning Mao et al. (2019).
@@
With Concepts spanning Neuroscience literature to interpretability literature, there are vast representations of concepts varying from concept activation vectors, concept trees, and graphs to concept activation areas in latent space. While some of the methods require concept definitions to be given by the user, several works automate this by automatically extracting concepts from models Ghorbani et al. (2019); Wang et al. (2022).
@@
Additionally, there are some recent methods using concepts for improving model performance by encoding prior knowledge or intervening and correcting the model to ensure it uses correct concepts. These methods are known as Concept Oriented Deep Learning (CODL) methods Sawada and Nakamura (2022b, a). We provide a systematic survey knitting this vast literature involving concepts.
@@ 2 Related work
Hitzler and Sarker (2022) present a survey on concept-based explainability techniques but are restricted to Concept Activation Vector (CAV) Kim et al. (2018) representations. Schwalbe (2022) survey various visual Concept Embeddings giving a taxonomy of Concept Analysis Approaches and various datasets used for Concepts Supervision. A parallel work Holmberg et al. (2022) also surveys knowledge representations mapping to concepts and classifies concept-based XAI into Associations (what if I see), Interventions (what if I do), and Counterfactuals (what if I had done). Gao et al. (2022a); Weber et al. (2022); Hase and Bansal (2021) survey various XAI-based approaches for model improvement. We categorize XAI-based model improvement in terms of the end goal as (i) Enhanced interpretability and (ii) Better Generalization. Most of the existing literature focuses on the former (i.e., enhancing interpretability of models) Koh et al. (2020); Kazhdan et al. (2020); Alvarez Melis and Jaakkola (2018); Teso (2019); Sawada and Nakamura (2022a, b); Stammer et al. (2021); Lage and Doshi-Velez (2020); Stammer et al. (2022); Teso et al. (2022); Friedrich et al. (2022); Keswani et al. (2022); Xue et al. (2022); Wang et al. (2023); Sacha et al. (2023) while some works address the latter (using interpretability to aid more generalizable models) Friedrich et al. (2022); Anders et al. (2022); Bahadori and Heckerman (2020); Stammer et al. (2021); Song et al. (2023); Bontempelli et al. (2023); Shao et al. (2022); Shu et al. (2019); Schramowski et al. (2020); Kronenberger and Haselhoff (2020); Bontempelli et al. (2021). Such division concurs with Holmberg et al. (2022), who found that explanations with better understanding significantly differ from explanations with the goal of finding actionable decisions. Gao et al. (2022a) classify XAI-based model improvement in terms of the type of explanation method used while Weber et al. (2022) classify based on augmentation made for improvement, which can be over data, loss, gradient, model, etc. Explanation Guided Learning (EGL) involves using explanations for model improvement to meet one or both of the above-listed goals (i, ii). Various local EGL techniques are surveyed by Gao et al. (2022a). Explanatory Interactive Learning (XIL) involves interactively using explanations with human feedback to improve the model. Friedrich et al. (2022) surveys various XIL techniques using local XAI methods to remove shortcuts in the models. eXplanatory Active Learning (XAL) involves using Explanations in an Active Learning framework as surveyed by Ghai et al. (2021).
@@
Hase and Bansal (2021) list papers of XAI in Natural Language Processing doing model improvement while we focus on Computer Vision XAI, specifically visual concepts-based model improvement. Beckh et al. (2021) surveys methods for model improvement for better explainability (type ii) using Prior knowledge. There is another line of work over CODL Chang (2018a), which provides a framework for Concept Oriented Deep Learning via Concept Graphs Chang (2018b). On the other hand, we provide a novel taxonomy of concept representations and a categorization of various Concept Discovery Algorithms and Concept Oriented Deep Learning (CODL). We restrict ourselves to visual representations of concepts and show applications of CODL in vision. We organize our survey based on representations for Concepts used by the existing works followed by concept discovery methods, concept-based model improvement, evaluation of discovered concepts, and evaluation by concepts.
@@ 3 Various Concept Representations
The landscape of model interpretability, particularly through the lens of concept-based explanations, is rich and varied. This section navigates through the myriad of concept representations, elucidating their unique contributions and applications in deep learning models.
@@
Table 2: Concept Discovery categorization
Type Method used Papers
Post-hoc Super-pixel
ACE Ghorbani et al. (2019), CocoX Akula et al. (2020)
Coorperative Attribution (shapeley)
ConceptSHAP Yeh et al. (2020)
Autoencoders based
PACE Kamakshi et al. (2021)
Causality based
MCE Vielhaben et al. (2023), Alipour et al. (2022)
Using Probes
ACDTE Shawi et al. (2021), Silver et al. (2017)
Using Convex Optimization
Schut et al. (2023)
Ante-hoc Saliency based
HINT Wang et al. (2022), Kori et al. (2022b)
Using Slot-attention Locatello et al. (2020)
Stammer et al. (2021)
Latent Space Disentanglement based
XIL based
iCSNs Stammer et al. (2022), RRC Stammer et al. (2021)
@@ 7.1.1 Aiming better interpretability
      $\hat{q}$ is approximated in a supervised manner using a mapped set of concept labels and target labels (class labels)as a training set and fitting using Decision trees or logistic regression. Concept Bottleneck Model with Additional Unsupervised Concepts (CBM-AUC) Sawada and Nakamura (2022b) combines CBMs with SENNs to extend CBMs to additional unsupervised concepts along with supervised concepts. Grupen et al. (2022) introduce Concept Bottleneck Policies (CBPs) for enhancing interpretability in multi-agent reinforcement learning (MARL). CBPs use conditioning similar to CBMs but in action space in RL instead of classification space in CBMs. They first predict concepts given a state, then use the predicted concepts to select actions, thus enhancing interpretability. Das et al. (2023) presents a novel approach to generating concept-based explanations in sequential decision-making settings, focusing on both improving AI agents’ learning rates and enhancing end-user comprehension of AI decisions. They learn a joint embedding model to map state-action pairs to concept-based explanations. These explanations are then used for two primary purposes: informing reward shaping during the agent’s training and providing end-users with understandable insights into the agent’s decision-making process at deployment.
@@ 10 Conclusion
We thus discussed various aspects of concept based approaches starting from various methods in concept representation methods used in the interpretation of trained models to various concept discovery methods. We provided a detailed hierarchy of the methods and also discussed the concept evaluation metrics, followed by an explanation of the guided learning hierarchy, which focused on concept-oriented deep learning (CODL) methods. We classified various concept-based model improvement methods on the basis of two aims: better interpretability and better generalization. We further categorized them by discussing various existing methods. We also discussed applications of concept-guided methods in various fields like domain adaptation, AlphaZero games, and applications in medical domains. We start with post-hoc concept-based interpretability methods, discuss ante-hoc concept-based training methods, and finally discuss Concept-based Model Performance Evaluation.

### 23 Mar 2024 v2

@@ Abstract
The focus of recent research has shifted from merely improving the metrics based performance of Deep Neural Networks (DNNs) to DNNs which are more interpretable to humans. The field of eXplainable Artificial Intelligence (XAI) has observed various techniques, including saliency-based and concept-based approaches. These approaches explain the model’s decisions in simple human understandable terms called Concepts. Concepts are known to be the thinking ground of humans. Explanations in terms of concepts enable detecting spurious correlations, inherent biases, or clever-hans. With the advent of concept-based explanations, a range of concept representation methods and automatic concept discovery algorithms have been introduced. Some recent works also use concepts for model improvement in terms of interpretability and generalization. We provide a systematic review and taxonomy of various concept representations and their discovery algorithms in DNNs, specifically in vision. We also provide details on concept-based model improvement literature marking the first comprehensive survey of these methods.
@@ 1 Introduction
With the increasing use of deep learning in various tasks, there is rapid growth in the XAI literature. One of the most promising methods of XAI is concept-based approaches, which explain model predictions in human-understandable units called Concepts Kazhdan et al. (2021). Concepts are linked to their origin in NeuroScience Literature, in which they are defined as ideas derived or inferred from facts Donham (2010) and are known to hold the human’s mental world together Murphy (2002). It has been observed that humans learn by creating concept models of things around them, helping them gather generalized knowledge Lake et al. (2017). There are various spurious correlations which creep in DNNs due to their black box nature. Model explanations in terms of concepts not only lead to better interpretability but also help to identify these spurious correlations and inherent biases also known as Clever-hans.
@@
The interpretability research aims to explain the model’s decisions in simple human understandable terms Zhang et al. (2021b). Concepts, by their definition, are humanly understandable and, hence, a good choice for interpretability. Recent XAI literature has observed a surge in concept-oriented explanations. There have been many methods that derive concept-based explanations from pre-trained models (post-hoc), while some do it while training the model (ante-hoc). Kim et al. (2018) paved the way for post-hoc concept-based explanations of models using Concept Activation Vectors, while a parallel work Fong and Vedaldi (2018) explained Convolutional Neural Networks (CNN’s) filters in terms of concepts. Many works Bai et al. (2022); Chen et al. (2020); Chang (2018b); Bontempelli et al. (2023) build on the above concept representations in a post-hoc manner. Concept Bottleneck Models (CBMs) Koh et al. (2020) derive concept explanations ante-hoc, putting the weightage of concept explanations on the model while training. Recent work by Yuksekgonul et al. (2022) extended CBMs to post-hoc explanations.
@@
Concepts are essential for both human-like learning and better generalization by models. A natural intuition for human-like learning is making the model match the human attention for several tasks as done by Shen et al. (2021); Gao et al. (2022b); Liao et al. (2022); He et al. (2022); Helgstrand and Hultin (2022). One major limitation of such approaches is the requirement of additional human-level attention data. Chang (2018a) formalize Concept Oriented Deep Learning (CODL) for human alike learning involving concepts. They argue that CODL overcomes several limitations of deep learning, involving concerns of interpretability, transferability, contextual adaptation, and humongous training data requirements. For the scope of this work, we focus on concept-based literature and refrain from providing generic definitions of interpretability/explainability (or XAI) literature, referring the readers to Zhang et al. (2021b); Du et al. (2019); Linardatos et al. (2020); Vojíř and Kliegr (2020); Akhtar (2023) which present surveys on XAI methods. Another concept-inspired learning paradigm is Neuro-Symbolic learning, which involves learning like humans using symbolic reasoning Mao et al. (2019).
@@
With Concepts spanning Neuroscience literature to interpretability literature, there are vast representations of concepts varying from concept activation vectors, concept trees, and graphs to concept activation areas in latent space. While some of the methods require concept definitions to be given by the user, several works automate this by extracting concepts from models Ghorbani et al. (2019); Wang et al. (2022).
@@
Additionally, some recent methods use concepts for improving model performance by encoding prior knowledge or intervening and correcting the model to ensure it uses correct concepts Sawada and Nakamura (2022b, a). We provide a systematic survey knitting this vast literature involving concepts.
@@ 2 Related work
Hitzler and Sarker (2022) present a survey on concept-based explainability techniques but are restricted to Concept Activation Vector (CAV) Kim et al. (2018) representations. Schwalbe (2022) survey various visual concept embeddings giving a taxonomy of concept analysis approaches and various datasets used for supervision of concept learning. A parallel work Holmberg et al. (2022) also surveys knowledge representations mapping to concepts and classifies concept-based XAI into Associations (what if I see), Interventions (what if I do), and Counterfactuals (what if I had done).
@@
Gao et al. (2022a); Weber et al. (2022); Hase and Bansal (2021) survey various XAI-based approaches for model improvement. We categorize XAI-based model improvement in terms of the end goal as (i) Enhanced interpretability and (ii) Better Generalization. Most of the existing literature focuses on the former (i.e., enhancing interpretability of models) Koh et al. (2020); Kazhdan et al. (2020); Alvarez Melis and Jaakkola (2018); Teso (2019); Sawada and Nakamura (2022a, b); Stammer et al. (2021); Lage and Doshi-Velez (2020); Stammer et al. (2022); Teso et al. (2022); Friedrich et al. (2022); Keswani et al. (2022); Xue et al. (2022); Wang et al. (2023); Sacha et al. (2023) while some works address the latter (using interpretability to aid more generalizable models) Friedrich et al. (2022); Anders et al. (2022); Bahadori and Heckerman (2020); Stammer et al. (2021); Song et al. (2023); Bontempelli et al. (2023); Shao et al. (2022); Shu et al. (2019); Schramowski et al. (2020); Kronenberger and Haselhoff (2020); Bontempelli et al. (2021). Such division concurs with Holmberg et al. (2022), who found that explanations with better understanding significantly differ from explanations with the goal of finding actionable decisions. Gao et al. (2022a) classify XAI-based model improvement in terms of the type of explanation method used while Weber et al. (2022) classify based on augmentation made for improvement, which can be over data, loss, gradient, model, etc. Explanation Guided Learning (EGL) involves using explanations for model improvement to meet one or both of the above-listed goals (i, ii). Various local EGL techniques are surveyed by Gao et al. (2022a). Explanatory Interactive Learning (XIL) involves interactively using explanations with human feedback to improve the model. Friedrich et al. (2022) survey various XIL techniques using local XAI methods to remove shortcuts in the models. eXplanatory Active Learning (XAL) involves using Explanations in an Active Learning framework as surveyed by Ghai et al. (2021).
@@
Hase and Bansal (2021) survey XAI based model improvement methods in Natural Language Processing while we focus on such methods in Computer Vision, specifically visual concepts-based model improvement. Beckh et al. (2021) survey model improvement methods aiming better explainability (type ii) using Prior knowledge. Chang (2018b) provide framework for CODL Chang (2018a) using Concept Graphs. On the other hand, we provide a novel taxonomy of concept representations and a categorization of various Concept Discovery Algorithms and CODL. We restrict ourselves to visual representations of concepts and show applications of CODL in vision. We organize our survey based on representations for Concepts used by the existing works followed by concept discovery methods, concept-based model improvement, evaluation of discovered concepts, and evaluation by concepts.
@@ 3 Various Concept Representations
The landscape of model interpretability via the lens of concept-based explanations, is rich and varied. This section navigates through various concept representations methods highlighting their unique contributions and applications in deep learning models.
@@ 4 Concept Discovery Methods
Table 2: Concept Discovery categorization
Type Method used Papers
Post-hoc Super-pixel
ACE Ghorbani et al. (2019), CocoX Akula et al. (2020)
Coorperative Attribution (shapeley)
ConceptSHAP Yeh et al. (2020)
Autoencoders based
PACE Kamakshi et al. (2021)
Causality based
MCE Vielhaben et al. (2023), Alipour et al. (2022)
Using Probes
ACDTE Shawi et al. (2021), Silver et al. (2017)
Using Convex Optimization
Schut et al. (2023)
Ante-hoc Saliency based
HINT Wang et al. (2022), Kori et al. (2022b)
Using Slot-attention Locatello et al. (2020)
Stammer et al. (2021)
Latent Space Disentanglement based
XIL based
iCSNs Stammer et al. (2022), RRC Stammer et al. (2021)
@@ 7.1.1 Aiming better interpretability
      $\hat{q}$ is approximated in a supervised manner using a mapped set of concept labels and target labels (class labels)as a training set and fitting using Decision trees or logistic regression. Concept Bottleneck Model with Additional Unsupervised Concepts (CBM-AUC) Sawada and Nakamura (2022b) combines CBMs with SENNs to extend CBMs to additional unsupervised concepts along with supervised concepts. Grupen et al. (2022) introduce Concept Bottleneck Policies (CBPs) for enhancing interpretability in multi-agent reinforcement learning (MARL). CBPs use conditioning similar to CBMs but in action space in RL instead of classification space in CBMs. They first predict concepts given a state, then use the predicted concepts to select actions, thus enhancing interpretability. Das et al. (2023) present a novel approach to generating concept-based explanations in sequential decision-making settings, focusing on both improving AI agents’ learning rates and enhancing end-user comprehension of AI decisions. They learn a joint embedding model to map state-action pairs to concept-based explanations. These explanations are then used for two primary purposes: informing reward shaping during the agent’s training and providing end-users with understandable insights into the agent’s decision-making process at deployment.
@@ 10 Conclusion
We discussed various aspects of concept based approaches starting from concept representation methods to various concept discovery methods. We provided a detailed hierarchy of the methods and also discussed the concept evaluation metrics, followed by an explanation of the guided learning hierarchy, focusing on Concept-Oriented Deep Learning (CODL) methods. We classified concept-based model improvement methods on the basis of two aims: better interpretability and better generalization. We also discussed applications of concept-guided methods in fields like domain adaptation, AlphaZero games, and applications in medical domains. In summary, our discussion commenced from post-hoc concept-based interpretability methods, covered ante-hoc concept-based training methods and concluded with concept-based model performance evaluation.